# Defeaturing.bib

### v1

@@ Abstract
An adaptive refinement strategy, based on an equilibrated flux a posteriori error estimator, is proposed in the context of defeaturing problems. Defeaturing consists in removing features from complex domains in order to ease the meshing process, and to reduce the computational burden of simulations. It is a common procedure, for example, in computer aided design for simulation based manufacturing. However, depending on the problem at hand, the effect of geometrical simplification on the accuracy of the solution may be detrimental. The proposed adaptive strategy is hence twofold: starting from a defeatured geometry it allows both for standard mesh refinement and geometrical refinement, which consists in choosing, at each step, which features need to be included into the geometry in order to significantly increase the accuracy of the solution. With respect to other estimators that were previously proposed in the context of defeaturing, the use of an equilibrated flux reconstruction allows us to avoid the evaluation of the numerical flux on the boundary of features. This makes the estimator and the adaptive strategy particularly well-suited for finite element discretizations, in which the numerical flux is typically discontinuous across element edges. The inclusion of the features during the adaptive process is tackled by a CutFEM strategy, in order to preserve the non conformity of the mesh to the feature boundary and never remesh the computational domain as the features are added. Hence, the estimator also accounts for the error introduced by weakly imposing the boundary conditions on the boundary of the added features.
@@ 1 Introduction
Defeaturing consists in the simplification of a geometry by removing features that are considered not relevant for the approximation of the solution of a given PDE. It is a fundamental process to reduce the computational effort when repeated simulations on complex geometries are required, in particular when the computational domain is characterized by the presence of features of different scales and shapes. This is the case, for example, of simulation-based manufacturing. However, identifying which features actually have a negligible impact on the solution may be not trivial. Historically, the problem has been approached by exploiting some a priori knowledge of the domain, of the materials and of the problem at hand ([FINE00, FOUCAULT2004, THAKUR2009]). Nevertheless, an a posteriori estimator becomes necessary when the industrial design process needs to be automatized, and several examples can be found in the literature. In [Ferrandes2009] the defeaturing error is assumed to be concentrated on the feature boundary and feature-local problems are solved in order to estimate it. Alternative approaches are instead proposed in [CHOI2005, GOP2007, GOP2009, SOK1999, TUR2009], resorting to the concept of feature sensitivity analysis. In [LI2011b, LI2011, LI2013, LI2013b] an a posteriori estimator is built by reformulating the defeaturing error as a modeling error, whereas [TANG2013] resorts to the reciprocal theorem, stating flux conservation in the features.
@@
Recently, a precise framework for analysis-aware defeaturing for the Poisson equation was introduced in [BCV2022], proposing an estimator which explicitly depends on the feature size. A geometric adaptive strategy based on such an estimator was devised in [AC_2023], allowing one to choose, at each step, which features have a significant impact on the accuracy of the solution and should hence be included into the geometry. In [BCV2022_arxiv], geometric adaptivity was combined with standard mesh adaptivity, in an IGA discretization framework. These works were the starting point for the investigation carried out in [BCGVV], in which the defeaturing problem was instead tackled by means of an equilibrated flux a posteriori error estimator, specifically designed to be used along with a standard finite element discretization. The work in [BCGVV] was intended as a preliminary analysis for the use of the equilibrated flux estimator in an adaptive strategy similar to the one devised in [BCV2022_arxiv], i.e. a twofold adaptive approach, accounting both for standard mesh refinement and geometric refinement. Until no feature is included, the equilibrated flux a posteriori error estimator can easily be computed on a mesh which is completely blind to the boundary of the features themselves. The equilibrated flux reconstruction is indeed built following the local equilibration procedure proposed in [BraessScho2008, EV2015], and the integrals on the feature boundary, which are required to evaluate the defeaturing component of the estimator, can always be computed, regardless of the intersections with the mesh elements. However, the procedure adopted in [BCGVV] to reconstruct the equilibrated flux is not designed for trimmed elements, and a remeshing of the computational domain would hence be needed each time a feature is included. From a computational cost perspective, this if of course not a feasible option.
In this work we hence propose the adaptation of the estimator designed in [BCGVV] to the case of trimmed geometries. In particular, we focus on the negative feature case, examining perforated domains in which the features are the holes themselves. We consider Neumann boundary conditions on the feature boundary and we adapt to our purpose the CutFEM strategy proposed in [puppi2021cut] in order to weakly enforce the boundary condition when computing the local flux reconstruction close to the included features. This results in an a posteriori error estimator bounding both the defeaturing component of the error, i.e. the error related to the features that are not (yet) included into the geometry, and the numerical component of the error. This second component includes the standard numerical error on elements that are not cut by an included feature, for which the difference between the reconstructed flux and the numerical flux provides an upper bound with unitary reliability constant ([AINSWORTH1997, DEST_MET1999, BraessScho2008, Luce_Wohlmuth]), and the error related to the weak imposition of the Neumann boundary conditions and to the consequent degradation of the mass balance constraint on cut elements.
@@ 2 Notation and model problem
Let $\Omega\subset\mathbb{R}^{d}$ be an open domain with boundary $\partial\Omega$, characterized by a finite set of open features $\mathcal{F}$, i.e. by some geometrical details of smaller scale. We denote by $F$ a generic feature in $\mathcal{F}$ and by $\partial F$ its boundary. For the sake of simplicity in the analysis that follows, we assume that both $\Omega$ and each $F\in\mathcal{F}$ are Lipschitz domains. Adopting the notation introduced in [BCV2022], a feature $F$ can, in general, be either negative or positive. We say that $F$ is a negative feature if $(\overline{F}\cap\overline{\Omega})\subset\partial\Omega$, whereas it is positive if $F\subset\Omega$ In practice, a negative feature corresponds to a portion of the domain in which some material has been removed, internally or at the boundary, while in the positive feature case some material has been added at the boundary (see Figure 1-left).
Figure 1: Left, domain with positive and negative features; right, corresponding defeatured geometry.
The simplified geometry obtained by neglecting all the features is called defeatured geometry and denoted by $\Omega_{0}$. It is obtained by cutting off all the positive features and by filling with material the negative ones (see Figure 1-right). For simplicity, we assume it to be a Lipschitz domain as well. In the present work we decide to focus on the negative feature case, hence working with perforated materials. For this reason, the defeatured geometry is defined as
$\Omega_{0}:=\text{int}\big{(}\overline{\Omega}\cup\{\overline{F},~{}F\in\mathcal{F}\}\big{)}$ (2)
For some observations on the extension to the positive feature case, we refer to Remark 3.
Let $\partial\Omega=\Gamma_{\mathrm{D}}\cup\Gamma_{\mathrm{N}}$, with $\Gamma_{\mathrm{D}}\cap\Gamma_{\mathrm{N}}=\emptyset$ and $\Gamma_{D}\neq\emptyset$, and, as in [BCGVV], we assume that $\partial F\cap\Gamma_{D}=\emptyset$, $\forall F\in\mathcal{F}$, which means that all the features lie on a Neumann boundary. Finally, let
@@
As in [BCV2022_arxiv], we do the following separability assumption:
@@
  • For every $F,F^{\prime}\in\mathcal{F}$, $F\neq F^{\prime}$, $\overline{F}\cap\overline{F^{\prime}}=\emptyset$.
@@
      the maximum number of superposed subdomains is limited and notably smaller than the total number of features.
As observed in [AC_2023], such an assumption is actually pretty weak: the first condition is easily satisfied since, if two features intersect each other they can simply be redefined as a single feature; the second condition allows for features that are arbitrarily close to one another, provided that the number of close features is bounded.
@@
Let $\mathcal{T}_{h}^{0}$ be a non-degenerate simplicial mesh covering $\Omega_{0}$, such that $\overline{\Omega_{0}}=\bigcup_{K\in\mathcal{T}_{h}^{0}}\overline{K}$. Hereby, we suppose that the mesh faces match with the boundaries $\Gamma_{\mathrm{D}}$, $\Gamma_{\mathrm{N}}^{0}:=(\Gamma_{\mathrm{N}}\setminus\gamma)\cup\gamma_{0}$, but we are not asking $\mathcal{T}_{h}^{0}$ to be conforming to the boundaries in $\gamma$. Let us introduce the space
$Q_{h}(\Omega_{0})=\mathcal{P}_{1}(\mathcal{T}_{h}^{0}):=\left\{q_{h}\in L^{2}(\Omega_{0}):~{}{q_{h}}|_{{K}}\in\mathcal{P}_{1}(K),~{}\forall K\in\mathcal{T}_{h}^{0}\right\},$ (7)
with $\mathcal{P}_{1}(K)$ denoting the space of polynomials of degree at most 1 on $K\in\mathcal{T}_{h}^{0}$, and let
@@
Let us split the set of the features in two disjoint subsets $\tilde{\mathcal{F}}$ and $\mathcal{F}^{\star}$, such that $\mathcal{F}=\tilde{\mathcal{F}}\cup\mathcal{F}^{\star}$ and $\tilde{\mathcal{F}}\cap\mathcal{F}^{\star}=\emptyset$, and let us define a partially defeatured geometry as
${\Omega^{\star}}:=\text{int}\big{(}\overline{\Omega}\cup\{\overline{F},~{}F\in\tilde{\mathcal{F}}\}\big{)}$ (10)
@@
$Q_{h}(\Omega_{{}_{\mathcal{T}}})=\mathcal{P}_{1}(\mathcal{A}_{h}):=\left\{q_{h}\in L^{2}(\Omega_{{}_{\mathcal{T}}}):~{}{q_{h}}|_{{K}}\in\mathcal{P}_{1}(K),~{}\forall K\in\mathcal{A}_{h}\right\},$ (15)
@@
$V_{h}^{0}(\Omega_{{}_{\mathcal{T}}})=\{v\in C(\overline{\Omega_{{}_{\mathcal{T}}}})\cap Q_{h}(\Omega_{{}_{\mathcal{T}}}):{v}|_{{\Gamma_{\mathrm{D}}}}=0\},$
$V_{h}(\Omega_{{}_{\mathcal{T}}})=\{v\in C(\overline{\Omega_{{}_{\mathcal{T}}}})\cap Q_{h}(\Omega_{{}_{\mathcal{T}}}):{v}|_{{\Gamma_{\mathrm{D}}}}=g_{\mathrm{D}}\},$
@@ 3 Error estimator via equilibrated flux reconstruction
Our aim is to control the energy norm of the error $||\nabla(u-u_{h}^{\star})||_{\Omega}$, where $u$ is the solution of (4) and $u_{h}^{\star}$ is the solution of (16). However, we want to achieve this without solving Problem (4). For this reason we propose an a posteriori error estimator defined uniquely on $\Omega^{\star}$ and based on an equilibrated flux reconstructed from $u_{h}^{\star}$. This estimator allows both for standard mesh refinement and for geometric refinement, which consists in choosing, at each step, which features need to be included into the geometry in order to significantly increase the accuracy of the solution. It is directly adapted from [BCGVV] in order to handle the case of trimmed elements. Indeed, starting from the totally defeatured geometry $\Omega_{0}$, we aim at gradually including new features without remeshing the domain, but only performing local refinements of the initial non-conforming mesh $\mathcal{T}_{h}^{0}$.
@@
Then, if $u^{\star}\in H^{1}_{g_{\mathrm{D}},\Gamma_{\mathrm{D}}}(\Omega^{\star})$ is the solution of (14), the theorem of Prager and Synge (see [Braess, PragerSynge]) states that, for $v\in H^{1}_{g_{\mathrm{D}},\Gamma_{\mathrm{D}}}({\Omega^{\star}})$,
$||\nabla(u^{\star}-v)||_{\Omega^{\star}}^{2}+||\nabla u^{\star}+\bm{\sigma}||_{\Omega^{\star}}^{2}=||\nabla v+\bm{\sigma}||_{\Omega^{\star}}^{2}.$ (18)
@@
$||\nabla(u^{\star}-u_{h}^{\star})||_{\Omega^{\star}}^{2}\leq||\nabla u_{h}^{\star}+\bm{\sigma}||_{\Omega^{\star}}^{2},$
@@
Equilibrated fluxes are about the cheap construction of such a $\bm{\sigma}$ at a discrete level. In order to build $\bm{\sigma}_{h}$, we resort to a local equilibration procedure directly adapted from [BraessScho2008, EV2015]. Under this approach, local equilibrated fluxes are built on patches of elements sharing a node, avoiding to solve a global optimization problem and making the method well-suited for parallel implementation.
However, some extra care is needed in order to account for the non-conformity of the mesh with respect to ${\gamma^{\star}}$. Indeed, the Neumann boundary condition is essential for the flux, and it can be imposed strongly on $\Gamma_{\mathrm{N}}^{\star}$ in the definition of the discrete space, since we have assumed that the mesh is conforming to this portion of the boundary. The Neumann boundary condition on ${\gamma^{\star}}$ has instead to be imposed weakly: we hence decide to work in a CutFEM framework, adopting a Nitsche method as proposed in [puppi2021cut] for the case of the mixed formulation of the Darcy problem in the unfitted case. For a flux recovery for CutFEM in the case of Dirichlet boundary conditions we refer the reader to [HeCapatina], which resorts however to a different local equilibration procedure.
@@
In the following, we denote by ${\mathscr{N}_{h}}$ and $\mathscr{E}_{h}$ respectively the set of vertices and edges in the active mesh ${\mathcal{A}_{h}}$. In particular the set of vertices is decomposed into $\mathscr{N}_{h}^{\mathrm{ext}}$ and $\mathscr{N}_{h}^{\mathrm{int}}$, corresponding respectively to vertices lying on $\Gamma_{\mathrm{N}}^{\star}$ and inside $\Omega_{0}$. Let us remark that the nodes belonging to $\mathcal{A}_{h}$ but lying inside a feature will still belong to $\mathscr{N}_{h}^{\mathrm{int}}$.
@@
$\bm{M}_{h}({\omega_{\bm{a}}}):=\left\{\bm{v}_{h}\in\bm{H}(\text{div},{\omega_{\bm{a}}}):~{}{\bm{v}_{h}}|_{{K}}\in[\mathcal{P}_{1}(K)]^{d}+\bm{x}\mathcal{P}_{1}(K),~{}\forall K\in{\omega_{\bm{a}}}\right\},$
which is the Raviart-Thomas finite element space of order 1 on ${\omega_{\bm{a}}}$, and let
@@
$\bm{M}_{h}^{\bm{a}}$ :=$:=\begin{cases}\bm{M}_{h}^{\bm{a},0}&\text{ if }\bm{a}\in\mathscr{N}_{h}^{\mathrm{int}}\\ \big{\{}\bm{v}_{h}\in\bm{M}_{h}({\omega_{\bm{a}}}):~{}\bm{v}_{h}\cdot\bm{n}_{{\omega_{\bm{a}}}}=0\text{ on }{\partial\omega_{\bm{a}}^{0}}\\ \hskip 76.82234pt\bm{v}_{h}\cdot\bm{n}_{{\omega_{\bm{a}}}}=-{\psi_{\bm{a}}}g_{\mathrm{N}}\text{ on }{\partial\omega_{\bm{a}}^{\psi}}\cap\Gamma_{\mathrm{N}}^{0}\}&\text{ if }\bm{a}\in{\mathscr{N}_{h}}^{\mathrm{ext}},\end{cases}$ (19)
$Q_{h}^{\bm{a}}$ :=$:=\begin{cases}\big{\{}q_{h}\in Q_{h}({\omega_{\bm{a}}}):(q_{h},1)_{{\omega_{\bm{a}}}}=0\big{\}}&\text{ if }\bm{a}\in{\mathscr{N}_{h}}^{\mathrm{int}}\text{ or }\bm{a}\in\text{int}(\overline{\Gamma_{\mathrm{N}}^{0}})\\[3.99994pt] Q_{h}({\omega_{\bm{a}}})&\text{ if }\bm{a}\in{\mathscr{N}_{h}}^{\mathrm{ext}}\text{ and }\bm{a}\notin\text{int}(\overline{\Gamma_{\mathrm{N}}^{0}}),\end{cases}$ (20)
@@
For $p\in\{0,1\}$, let us then introduce the bilinear forms $m_{\bm{a}}:\bm{M}_{h}^{\bm{a}}\times\bm{M}_{h}^{\bm{a}}\rightarrow\mathbb{R}$ and $b_{\bm{a}}^{p}:\bm{M}_{h}^{\bm{a}}\times Q_{h}^{\bm{a}}\rightarrow\mathbb{R}$ such that
$m_{\bm{a}}(\bm{\rho}_{h},\bm{v}_{h}):=(\bm{\rho}_{h},\bm{v}_{h})_{{\omega}_{\bm{a}}^{\star}}+\frac{1}{h_{\bm{a}}}\langle\bm{\rho}_{h}\cdot\bm{n},\bm{v}_{h}\cdot\bm{n}\rangle_{\gamma_{\bm{a}}^{\star}}$
$b_{\bm{a}}^{p}(\bm{v}_{h},q_{h}):=(q_{h},\nabla\cdot\bm{v}_{h})_{{\omega}_{\bm{a}}^{\star}}-p\langle q_{h},\bm{v}_{h}\cdot\bm{n}\rangle_{\gamma_{\bm{a}}^{\star}}$
and the linear operators $L_{\bm{a}}:\bm{M}_{h}^{\bm{a}}\rightarrow\mathbb{R}$ and $R_{\bm{a}}^{p}:Q_{h}^{\bm{a}}\rightarrow\mathbb{R}$ such that
@@
$R_{\bm{a}}^{p}(q_{h}):=(\psi_{\bm{a}}f-\nabla\psi_{\bm{a}}\cdot\nabla u_{h}^{\star},q_{h})_{{\omega}_{\bm{a}}^{\star}}+p\langle\psi_{\bm{a}}g,q_{h}\rangle_{\gamma_{\bm{a}}^{\star}},$ (22)
@@
$\begin{cases}m_{\bm{a}}(\bm{\sigma}_{h}^{\bm{a}},\bm{v}_{h})-b_{\bm{a}}^{1}(\bm{v}_{h},\lambda_{h}^{\bm{a}})=L_{\bm{a}}(\bm{v}_{h})&\forall\bm{v}_{h}\in\bm{M}_{h}^{\bm{a}}\\ \hskip 62.59596ptb_{\bm{a}}^{p}(\bm{\bm{\sigma}_{h}^{\bm{a}}},q_{h})=R_{\bm{a}}^{p}(q_{h})&\forall q_{h}\in Q_{h}^{\bm{a}},\end{cases}$ (23)

### v2

# Adaptive refinement in defeaturing problems via an equilibrated flux a posteriori error estimator

@@ Abstract
An adaptive refinement strategy, based on an equilibrated flux a posteriori error estimator, is proposed in the context of defeaturing problems. Defeaturing consists of removing features from complex domains to simplify mesh generation and reduce the computational cost of simulations. It is a common procedure, for example, in computer aided design for simulation-based manufacturing. However, depending on the problem at hand, geometrical simplification may significantly deteriorate the accuracy of the solution. The proposed adaptive strategy is hence twofold: starting from a defeatured geometry, it performs both standard mesh refinement and geometrical refinement by selecting, at each step, which features must be reintroduced to significantly improve accuracy. Similar adaptive strategies have been previously developed using residual-based error estimators within an IGA framework. Here, instead, we extend a previously developed equilibrated flux a posteriori error analysis, designed for standard finite element discretizations, to make it fully applicable within the adaptive procedure. In particular, we address the assembly of the equilibrated flux estimator in presence of elements trimmed by the boundary of included features, adopting a CutFEM strategy to handle feature inclusion. The resulting estimator allows us to bound both the defeaturing and the numerical sources of error, with additional contributions accounting for the weak imposition of boundary conditions.
@@ 1 Introduction
Defeaturing consists of the simplification of a geometry by removing features that are considered not relevant for the approximation of the solution of a given PDE. It is a fundamental process to reduce the computational effort when repeated simulations on complex geometries are required, in particular when the computational domain is characterized by the presence of features of different scales and shapes. This is the case, for example, of simulation-based manufacturing. However, identifying which features actually have a negligible impact on the solution may be not trivial. Historically, the problem has been approached by exploiting some a priori knowledge of the domain, of the materials and of the problem at hand ([19, 20, 37]). Nevertheless, an a posteriori estimator becomes necessary when the industrial design process needs to be automatized, and several examples can be found in the literature. In [18] the defeaturing error is assumed to be concentrated on the feature boundary and feature-local problems are solved in order to estimate it. Alternative approaches are instead proposed in [13, 22, 23, 35, 38], resorting to the concept of feature sensitivity analysis. In [28, 25, 27, 26] an a posteriori estimator is built by reformulating the defeaturing error as a modeling error, whereas [36] resorts to the reciprocal theorem, stating flux conservation in the features.
@@
Recently, a precise framework for analysis-aware defeaturing for the Poisson equation was introduced in [8], proposing an estimator which explicitly depends on the feature size. A geometric adaptive strategy based on this estimator was devised in [2], enabling, at each step, the selective inclusion of features according to their impact on the solution accuracy. In [10], geometric adaptivity was integrated with standard mesh adaptivity within an IGA discretization framework, by coupling the defeaturing estimator of [8] with a residual-based a posteriori estimator for the numerical error. These works were the starting point for the investigation carried out in [7], where the residual-based estimator was replaced by an equilibrated flux estimator, bounding the numerical component of the error with no unknown constant ([1, 15, 4, 29]). Employing the equilibrated flux also within the defeaturing component made the estimator particularly suitable for standard finite element discretizations, in which the numerical flux is typically discontinuous across element faces. In [7] the equilibrated flux is reconstructed on a mesh that is completely unaware of the presence of the features, using the local equilibration procedure proposed in [4, 17], and integrals on the feature boundary, which are required to evaluate the defeaturing component of the estimator, can always be computed, regardless of the intersections with the mesh elements.
The analysis presented in [7] was intended as a preliminary step towards the use of the equilibrated flux estimator in an adaptive strategy similar to the one devised in [10], combining mesh and geometric refinement. However, the a posteriori error analysis carried out in [7] remains valid only until the first feature is marked for inclusion. Once a feature is included, the mesh becomes non-conforming to an actual domain boundary, and the flux reconstruction and the estimator assembly need to be addressed in presence of trimmed mesh elements. Alternatively, repeated remeshing of the computational domain would be required after each feature inclusion, which is, of course, not a viable option from a computational cost perspective.
In this work we extend the analysis of [7] to build an a posteriori error estimator on partially defeatured geometries, where selected features have already been reintroduced into the domain. We focus on the negative Neumann feature case, namely small holes on which Neumann boundary conditions are imposed. Observing that the Neumann boundary condition is essential for the reconstructed flux, we adapt the CutFEM strategy proposed in [32] to weakly enforce the boundary condition on the immersed feature boundary during flux reconstruction. The resulting weak equilibration introduces additional contributions to the estimator. Nevertheless, the reconstruction is designed so that these additional terms vanish on elements that are not directly intersected by the feature boundary. The final a posteriori error estimator retains the decomposition into a defeaturing and a numerical component: the former bounds the error due to the features that are not yet included into the geometry, while the latter consists of the standard constant-free equilibrated flux estimator on uncut elements, together with additional contributions on trimmed elements. These contributions account for the weak imposition of Neumann boundary conditions and the resulting loss of local mass conservation.
@@ 2 Notation and model problem
Let $\Omega\subset\mathbb{R}^{d}$ be an open domain with boundary $\partial\Omega$, characterized by a finite set of open Neumann negative features $\mathcal{F}$. We say that $F\in\mathcal{F}$ is a negative feature of $\Omega$ if it is a geometric detail of smaller scale such that $(\overline{F}\cap\overline{\Omega})\subset\partial\Omega$. In practice, a negative feature corresponds to a portion of the domain in which some material has been removed, internally or at the boundary (see Figure 1-left). We call $F\in\mathcal{F}$ a Neumann feature if a Neumann boundary condition is imposed on $\partial F\cap\partial\Omega$. We refer to Remark 4 to some comments on different types of features and boundary conditions. For the sake of simplicity in the analysis that follows, we assume that both $\Omega$ and each $F\in\mathcal{F}$ are Lipschitz domains, and we denote by $\partial F$ the boundary of a generic feature $F$.
Figure 1: Left, domain with two negative features; right, corresponding defeatured geometry.
The simplified geometry obtained by neglecting all the features is called defeatured geometry. We denote it by $\Omega_{0}$ and, in practice, we obtain it by filling with material the negative features (see Figure 1-right), i.e.
$\Omega_{0}:=\text{int}\big(\overline{\Omega}\cup\{\overline{F},~F\in\mathcal{F}\}\big).$ (2)
For simplicity, we assume $\Omega_{0}$ to be a Lipschitz domain as well.
Let $\partial\Omega=\Gamma_{\mathrm{D}}\cup\Gamma_{\mathrm{N}}$, with $\Gamma_{\mathrm{D}}\cap\Gamma_{\mathrm{N}}=\emptyset$ and $\Gamma_{D}\neq\emptyset$, and, as in [7], we assume that $\partial F\cap\Gamma_{D}=\emptyset$, $\forall F\in\mathcal{F}$, which means that all the features lie on a Neumann boundary. Finally, let
@@
As in [10], we do the following separability assumption:
@@
  • For every $^{\prime}\in\mathcal{F}, FFF\neq F^{\prime}$, $\overline{F}\cap\overline{F^{\prime}}=\emptyset$.
@@
      the maximum number of superposed subdomains is bounded independently of the number of features.
As observed in [2], such an assumption is actually pretty weak: the first condition is easily satisfied since, if two features intersect each other they can simply be redefined as a single feature; the second condition allows for features that are arbitrarily close to one another, provided that the number of close features is bounded.
@@
Let $\mathcal{T}_{h}^{0}$ be a non-degenerate simplicial mesh covering $\Omega_{0}$, consisting of open simplices $K$, such that $\overline{\Omega_{0}}=\bigcup_{K\in\mathcal{T}_{h}^{0}}\overline{K}$. Hereby, we suppose that the mesh faces match with the boundaries $\Gamma_{\mathrm{D}}$, $\Gamma_{\mathrm{N}}^{0}:=(\Gamma_{\mathrm{N}}\setminus\gamma)\cup\gamma_{0}$, but we are not asking $\mathcal{T}_{h}^{0}$ to be conforming to the boundaries in $\gamma$. Let us introduce the space
$Q_{h}(\Omega_{0})=\mathcal{P}_{r}(\mathcal{T}_{h}^{0}):=\left\{q_{h}\in L^{2}(\Omega_{0}):~{q_{h}}|_{{K}}\in\mathcal{P}_{r}(K),~\forall K\in\mathcal{T}_{h}^{0}\right\},$ (7)
with $\mathcal{P}_{r}(K)$ denoting the space of polynomials of degree at most r{r} on $K\in\mathcal{T}_{h}^{0}$, with $r\geq 1$, and let
@@
Let us split the set of the features in two disjoint subsets $\tilde{\mathcal{F}}$ and $\mathcal{F}^{\star}$, such that $\mathcal{F}=\tilde{\mathcal{F}}\cup\mathcal{F}^{\star}$ and $\tilde{\mathcal{F}}\cap\mathcal{F}^{\star}$$\emptyset$, and let us define a partially defeatured geometry as
[…]
@@
$Q_{h}(\Omega_{{}_{\mathcal{T}}})=\mathcal{P}_{r}(\mathcal{A}_{h}):=\left\{q_{h}\in L^{2}(\Omega_{{}_{\mathcal{T}}}):~{q_{h}}|_{{K}}\in\mathcal{P}_{{r}}(K),~\forall K\in\mathcal{A}_{h}\right\},$ (15)
@@
$V_{h}^{0}(\Omega_{{}_{\mathcal{T}}})=\{v\in{\mathcal{C}^{0}}(\overline{\Omega_{{}_{\mathcal{T}}}})\cap Q_{h}(\Omega_{{}_{\mathcal{T}}}):{v}|_{{\Gamma_{\mathrm{D}}}}=0\},$
$V_{h}(\Omega_{{}_{\mathcal{T}}})=\{v\in{\mathcal{C}^{0}}(\overline{\Omega_{{}_{\mathcal{T}}}})\cap Q_{h}(\Omega_{{}_{\mathcal{T}}}):{v}|_{{\Gamma_{\mathrm{D}}}}=g_{\mathrm{D}}\},$
@@ 3 Error estimator via equilibrated flux reconstruction
Our aim is to control the energy norm of the error $\nabla(u-u_{h}^{\star})||_{\Omega}$, where $u$ is the solution of (4) and $u_{h}^{\star}$ is the solution of (16). However, we want to achieve this without solving Problem (4). For this reason we propose an a posteriori error estimator defined uniquely on $\Omega^{\star}$ and based on an equilibrated flux reconstructed from $u_{h}^{\star}$. This estimator allows both for standard mesh refinement and for geometric refinement, which consists in choosing, at each step, which features need to be included into the geometry in order to significantly increase the accuracy of the solution. It is directly adapted from [7] in order to handle the case of trimmed elements. Indeed, starting from the totally defeatured geometry $\Omega_{0}$, we aim at gradually including new features without remeshing the domain, but only performing local refinements of the initial non-conforming mesh $\mathcal{T}_{h}^{0}$.
@@
Then, if $u^{\star}\in H^{1}_{g_{\mathrm{D}},\Gamma_{\mathrm{D}}}(\Omega^{\star})$ is the solution of (14), the theorem of Prager and Synge (see [5, 31]) states that, for $v\in H^{1}_{g_{\mathrm{D}},\Gamma_{\mathrm{D}}}({\Omega^{\star}})$,
@@
Choosing $v=u_{h}^{\star}$ in (18) we obtain
[…]
@@
Equilibrated fluxes are about the cheap construction of such a $\bm{\sigma}$ at a discrete level. In order to build $\bm{\sigma}_{h}$, we resort to a local equilibration procedure directly adapted from [4, 17]. Under this approach, local equilibrated fluxes are built on patches of elements sharing a node, avoiding to solve a global optimization problem and making the method well-suited for parallel implementation.
However, some extra care is needed in order to account for the non-conformity of the mesh with respect to $\gamma^{\star}$}. Indeed, the Neumann boundary condition is essential for the flux, and it can be imposed strongly on $\Gamma_{\mathrm{N}}^{\star}$ in the definition of the discrete space, since we have assumed that the mesh is conforming to this portion of the boundary. The Neumann boundary condition on $\gamma^{\star}$} has instead to be imposed weakly: we hence decide to work in a CutFEM framework, adopting a Nitsche method as proposed in [32] for the case of the mixed formulation of the Darcy problem in the unfitted case. For a flux recovery for CutFEM in the case of Dirichlet boundary conditions we refer the reader to [14], which resorts however to a different local equilibration procedure.
@@
In the following, we denote by $\mathscr{N}_{h}$} and $\mathscr{E}_{h}$ respectively the set of vertices and faces (edges if $d=2$) in the active mesh $\mathcal{A}_{h}$}. In particular the set of vertices is decomposed into $\mathscr{N}_{h}^{\mathrm{ext}}$ and $\mathscr{N}_{h}^{\mathrm{int}}$, corresponding respectively to vertices lying on $\Gamma_{\mathrm{N}}^{\star}$ and inside $\Omega_{0}$. Let us remark that the nodes belonging to $\mathcal{A}_{h}$ but lying inside a feature will still belong to $\mathscr{N}_{h}^{\mathrm{int}}$.
@@
$\bm{M}_{h}({\omega_{\bm{a}}}):=\left\{\bm{v}_{h}\in\bm{H}(\text{div},{\omega_{\bm{a}}}):~{\bm{v}_{h}}|_{{K}}\in[\mathcal{P}_{r}(K)]^{d}+\bm{x}\mathcal{P}_{r}(K),~\forall K\in{\omega_{\bm{a}}}\right\},$
which is the Raviart-Thomas finite element space of order r{r} on $\omega_{\bm{a}}$}, and let
@@
[…]
@@
Let us then introduce the bilinear forms $m_{\bm{a}}:\bm{M}_{h}^{\bm{a}}\times\bm{M}_{h}^{\bm{a}}\rightarrow\mathbb{R}$ and ${b_{\bm{a}}}:\bm{M}_{h}^{\bm{a}}\times Q_{h}^{\bm{a}}\rightarrow\mathbb{R}$ such that
$m_{\bm{a}}(\bm{\rho}_{h},\bm{v}_{h}):=(\bm{\rho}_{h},\bm{v}_{h})_{{\omega}_{\bm{a}}^{\star}}+\frac{1}{h_{\bm{a}}}\langle\bm{\rho}_{h}\cdot\bm{n},\bm{v}_{h}\cdot\bm{n}\rangle_{\gamma_{\bm{a}}^{\star}}$ (21)
${b_{\bm{a}}}(\bm{v}_{h},q_{h}):=(q_{h},\nabla\cdot\bm{v}_{h})_{{\omega}_{\bm{a}}^{\star}}-{\langle q_{h},\bm{v}_{h}\cdot\bm{n}\rangle_{\gamma_{\bm{a}}^{\star}}}$ (22)
and the linear operators $L_{\bm{a}}:\bm{M}_{h}^{\bm{a}}\rightarrow\mathbb{R}$ and ${R_{\bm{a}}}:Q_{h}^{\bm{a}}\rightarrow\mathbb{R}$ such that
@@
${R_{\bm{a}}}(q_{h}):=(\psi_{\bm{a}}f-\nabla\psi_{\bm{a}}\cdot\nabla u_{h}^{\star},q_{h})_{{\omega}_{\bm{a}}^{\star}}+{\langle\psi_{\bm{a}}g,q_{h}\rangle_{\gamma_{\bm{a}}^{\star}}},$ (24)
@@
$\begin{cases}m_{\bm{a}}(\bm{\sigma}_{h}^{\bm{a}},\bm{v}_{h})-{b_{\bm{a}}}(\bm{v}_{h},\lambda_{h}^{\bm{a}})=L_{\bm{a}}(\bm{v}_{h})&\forall\bm{v}_{h}\in\bm{M}_{h}^{\bm{a}}\\ \hskip 62.59596pt{b_{\bm{a}}}(\bm{\bm{\sigma}_{h}^{\bm{a}}},q_{h})={R_{\bm{a}}}(q_{h})&\forall q_{h}\in Q_{h}^{\bm{a}},\end{cases}$ (25)